# Prot2Text: Multimodal Protein’s Function Generation with GNNs and Transformers

### v2

@@ Abstract
In recent years, significant progress has been made in this field of protein function prediction with the development of various machine-learning approaches. However, most existing methods formulate the task as a multi-classification problem, i.e. assigning predefined labels to proteins. In this work, we propose a novel approach, Prot2Text, which predicts a protein’s function in a free text style, moving beyond the conventional binary or categorical classifications. By combining Graph Neural Networks(GNNs) and Large Language Models(LLMs), in an encoder-decoder framework, our model effectively integrates diverse data types including protein sequence, structure, and textual annotation. This multimodal approach allows for a holistic representation of proteins’ functions, enabling the generation of detailed and accurate descriptions. To evaluate our model, we extracted a multimodal protein dataset from SwissProt, and demonstrate empirically the effectiveness of Prot2Text. These results highlight the transformative impact of multimodal models, specifically the fusion of GNNs and LLMs, empowering researchers with powerful tools for more accurate prediction of proteins’ functions.
@@ Results.
We report the results in Table 1, for different encoder models, including unimodal encoders like vanilla-Transformer, ESM2-35M, and RGCN, and multimodal encoders like RGCN vanilla-Transformer and RGCN ESM2-35. All models use a GPT-2 decoder. The unimodal vanilla-Transformer baseline, relying solely on the amino acid sequence of the protein, exhibits the lowest performance across all evaluation metrics. However, we observe a significant improvement in performance when using the unimodal graph encoder RGCN. The RGCN outperforms the vanilla-Transformer by over five absolute points in terms of BLEU score and three points in terms of BERT score. This performance disparity highlights the importance of incorporating structural information through the RGCN encoder for protein’s function prediction. On the other hand, leveraging the pretrained protein language model ESM2-35M instead of initializing the vanilla-Transformer randomly, results in a remarkable improvement in performance. The ESM2-35M encoder leads to a substantial increase of over 16 BLEU score points and 18 Rouge-L points compared to the standard vanilla-Transformer configuration. This notable enhancement can be attributed to the pretraining of ESM2-35M using masked protein modeling, which enables the encoder to capture intricate relationships and patterns within protein sequences. In the context of multimodal protein representation, the evaluation results demonstrate that Prot2Text exhibits superior performance across all assessment metrics. Notably, it achieves the highest BLEU Score of , the highest Rouge-1 score of , the highest Rouge-2 score of , the highest Rouge-L score of , and the highest BERT Score of . These outcomes highlight the effectiveness of fusing protein structure and amino acid information in a multimodal manner. The incorporation of protein structure, facilitated by the Relational Graph Convolutional Network (RGCN) with the sequential representations of amino acids from ESM2-35, significantly enhances the overall performance across all evaluation metrics. This improvement is attributed to the enriched understanding of proteins achieved through the synergy of these two modalities. Furthermore, the efficacy of the multimodal fusion approach is corroborated by the results obtained from RGCN vanilla-Transformer. Introducing structural information using RGCN to the randomly initialized vanilla-Transformer yields a substantial improvement of over BLEU score points compared to using the vanilla-Transformer alone, and more than BLEU score points improvement over using RGCN in isolation. Finally, to show the importance of the fusion block in the Prot2Text framework, we compare it against RGCN ESM2-25, which concatenates the protein structure representation to the amino acids representations. In this case, the graph representation will simply be passed to the decoder alongside the ESM output. We notice that using this strategy leads to slightly worse results than using the ESM alone. This not only provides backing for the selection of the fusion block employed in Prot2Text, but also suggests that indiscriminately increasing the overall parameter count of the model could potentially lead to a degradation in its performance.
@@
Ablation Study: Scaling to Larger Models.
We conducted an ablation study to assess the performance of our Prot2Text framework as we varied the number of parameters. The primary objective of this experiment was to evaluate the benefits of employing larger models in terms of generating more accurate and detailed textual representations of protein’s function. To conduct the ablation study, we systematically varied the size of the protein language model (ESM). Where Prot2Text, Prot2Text, Prot2Text and Prot2Text use ESM2-8M, ESM2-35M, ESM2-150M and ESM2-650M respectively. We evaluated each configuration on the same test set of proteins and used the same evaluation metrics as described earlier. The results of the ablation study, presented in Table 2, show a trend of performance improvement as we scale up the model’s architecture. Larger versions of ESM outperformed their smaller counterparts in most evaluation metrics. The increase in model size led to more accurate and relevant descriptions, indicating the benefit of leveraging larger language models in the Prot2Text framework. Yet, complementary analysis including corresponding computation time showed an increase in the inference cost following the use of larger models (higher number of parameters). Therefore, ( parameters) is a good trade-off striking the balance between performance and time cost. Furthermore, in Figure 2 we report the performance of all Prot2text models with respect to different similarity thresholds. Where the similarity represents the highest alignment score between the amino acid sequences of the test and train sets using BLAST identity. We observe that for test proteins with low similarity scores with the train set (between 20% and 30%) and for proteins with no counterpart in the train set, the Prot2Text is the dominant one while for higher similarity scores Prot2Text performs better.
@@ 5 Conclusion
In conclusion, our paper introduces Prot2Text, a pioneering multimodal framework, for the accurate prediction of a protein’s function in free text format, from graph and sequential input. By reformulating the task as free-text prediction, we address the limitations of traditional classification-based methods, allowing for a more nuanced and in-depth understanding of a protein’s functionality. Leveraging the power of GNNs and LLMs, we integrate structural and textual protein information, resulting in highly detailed and coherent generated protein descriptions. The release of a comprehensive multimodal protein dataset further empowers the scientific community to benchmark and advance the field of protein function prediction in free text format. This innovative approach opens new horizons for research and applications in drug discovery, protein engineering, and various biological sciences, with the potential to revolutionize our understanding of proteins’ functions.
@@ A.2 Text Generation
To generate the protein textual description during and after the training, we used the generation function implemented in the transformers library. We used the default generation parameters of length_penalty=1.0, no_repeat_ngram_size=0 and early_stopping=False. The text generation was done during the training on the validation set each 500 training steps using greedy search (number of beams equal to one) with maximum length of 256 tokens per sample. However, different configuration could be used leading to multiple functions. For example, the generated text in figure 3 from the main paper uses length_penalty=2.0, no_repeat_ngram_size=3 and early_stopping=True using Prot2Text.
Figure 5 shows the BLEU score validation throughout the training for the Prot2Text model. The validation BLEU score start to stabalize after the 20th epochs reaching the best validation BLEU score of 37.09 at the step 23000.
@@ A.3 CO2 Emission Related to Experiments
Experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.057 kgCOeq/kWh. A cumulative of 23000 hours of computation was performed on hardware of type Tesla V100-SXM2-32GB (TDP of 300W). Total emissions are estimated to be 393.3 kgCOeq of which 0 percents were directly offset. Estimations were conducted using the MachineLearning Impact calculator222https://mlco2.github.io/impact#compute presented in Lacoste et al., (2019).

### v3

@@ Abstract
In recent years, significant progress has been made in the field of protein function prediction with the development of various machine-learning approaches. However, most existing methods formulate the task as a multi-classification problem, i.e. assigning predefined labels to proteins. In this work, we propose a novel approach, Prot2Text, which predicts a protein’s function in a free text style, moving beyond the conventional binary or categorical classifications. By combining Graph Neural Networks(GNNs) and Large Language Models(LLMs), in an encoder-decoder framework, our model effectively integrates diverse data types including protein sequence, structure, and textual annotation and description. This multimodal approach allows for a holistic representation of proteins’ functions, enabling the generation of detailed and accurate functional descriptions. To evaluate our model, we extracted a multimodal protein dataset from SwissProt, and demonstrate empirically the effectiveness of Prot2Text. These results highlight the transformative impact of multimodal models, specifically the fusion of GNNs and LLMs, empowering researchers with powerful tools for more accurate function prediction of existing as well as first-to-see proteins.
@@ Results.
We report the results in Table 1, for different encoder models, including unimodal encoders like vanilla-Transformer, ESM2-35M, and RGCN, and multimodal encoders like RGCN vanilla-Transformer and RGCN ESM2-35. All models use a GPT-2 decoder. The unimodal vanilla-Transformer baseline, relying solely on the amino acid sequence of the protein, exhibits the lowest performance across all evaluation metrics. However, we observe a significant improvement in performance when using the unimodal graph encoder RGCN. The RGCN outperforms the vanilla-Transformer by over five absolute points in terms of BLEU score and three points in terms of BERT score. This performance disparity highlights the importance of incorporating structural information through the RGCN encoder for protein’s function prediction. On the other hand, leveraging the pretrained protein language model ESM2-35M instead of initializing the vanilla-Transformer randomly, results in a remarkable improvement in performance. The ESM2-35M encoder leads to a substantial increase of over 16 BLEU score points and 18 Rouge-L points compared to the standard vanilla-Transformer configuration. This notable enhancement can be attributed to the pretraining of ESM2-35M using masked protein modeling, which enables the encoder to capture intricate relationships and patterns within protein sequences. In the context of multimodal protein representation, the evaluation results demonstrate that Prot2TextBASE exhibits superior performance across all assessment metrics. Notably, it achieves the highest BLEU Score of , the highest Rouge-1 score of , the highest Rouge-2 score of , the highest Rouge-L score of , and the highest BERT Score of . These outcomes highlight the effectiveness of fusing protein structure and amino acid information in a multimodal manner. The incorporation of protein structure, facilitated by the Relational Graph Convolutional Network (RGCN) with the sequential representations of amino acids from ESM2-35, significantly enhances the overall performance across all evaluation metrics. This improvement is attributed to the enriched understanding of proteins achieved through the synergy of these two modalities. Furthermore, the efficacy of the multimodal fusion approach is corroborated by the results obtained from RGCN vanilla-Transformer. Introducing structural information using RGCN to the randomly initialized vanilla-Transformer yields a substantial improvement of over BLEU score points compared to using the vanilla-Transformer alone, and more than BLEU score points improvement over using RGCN in isolation. Finally, to show the importance of the fusion block in the Prot2Text framework, we compare it against RGCN ESM2-25, which concatenates the protein structure representation to the amino acids representations. In this case, the graph representation will simply be passed to the decoder alongside the ESM output. We notice that using this strategy leads to slightly worse results than using the ESM alone. This not only provides backing for the selection of the fusion block employed in Prot2Text, but also suggests that indiscriminately increasing the overall parameter count of the model could potentially lead to a degradation in its performance.
@@
Scaling to Larger Models.
We conducted an ablation study to assess the performance of our Prot2Text framework as we varied the number of parameters. The primary objective of this experiment was to evaluate the benefits of employing larger models in terms of generating more accurate and detailed textual representations of protein’s function. To conduct the ablation study, we systematically varied the size of the protein language model (ESM). Where Prot2TextSMALL, Prot2TextBASE, Prot2TextMEDIUM and Prot2TextLARGE use ESM2-8M, ESM2-35M, ESM2-150M and ESM2-650M respectively. We evaluated each configuration on the same test set of proteins and used the same evaluation metrics as described earlier. The results of the ablation study, presented in Table 2, show a trend of performance improvement as we scale up the model’s architecture. Larger versions of ESM outperformed their smaller counterparts in most evaluation metrics. The increase in model size led to more accurate and relevant descriptions, indicating the benefit of leveraging larger language models in the Prot2Text framework. Yet, complementary analysis including corresponding computation time showed an increase in the inference cost following the use of larger models (higher number of parameters). Therefore, ( parameters) is a good trade-off striking the balance between performance and time cost. Furthermore, in Figure 2 we report the performance of all Prot2text models with respect to different similarity thresholds. Where the similarity represents the highest alignment score between the amino acid sequences of the test and train sets using BLAST identity. We observe that for test proteins with low similarity scores with the train set (between 20% and 30%) and for proteins with no counterpart in the train set, the Prot2TextMEDIUM is the dominant one while for higher similarity scores Prot2TextLARGE performs better.
@@ 5 Conclusion
In conclusion, our paper introduces Prot2Text, a pioneering multimodal framework, for the accurate prediction of a protein’s function in free text format, from graph and sequential input. By reformulating the task as free-text prediction, we address the limitations of traditional classification-based methods, allowing for a more nuanced and in-depth understanding of a protein’s functionality. Leveraging the power of GNNs and LLMs, we integrate structural and textual protein information, resulting in highly detailed and coherent generated protein descriptions. The release of a comprehensive multimodal protein dataset further empowers the scientific community to benchmark and advance the field of protein function prediction in free text format. This innovative approach opens new horizons for research and applications in drug discovery, protein engineering, and various biological sciences, with the potential to revolutionize our understanding of proteins’ functions. This work is accepted at AAAI 2024222https://ojs.aaai.org/index.php/AAAI/article/view/28948.
@@ A.2 Text Generation
To generate the protein textual description during and after the training, we used the generation function implemented in the transformers library. We used the default generation parameters of length_penalty=1.0, no_repeat_ngram_size=0 and early_stopping=False. The text generation was done during the training on the validation set each 500 training steps using greedy search (number of beams equal to one) with maximum length of 256 tokens per sample. However, different configuration could be used leading to multiple functions. For example, the generated text in figure 3 uses length_penalty=2.0, no_repeat_ngram_size=3 and early_stopping=True using Prot2TextBASE. Figure 5 shows the BLEU score validation throughout the training for the Prot2TextBASE model. The validation BLEU score start to stabalize after the 20th epochs reaching the best validation BLEU score of 37.09 at the step 23000.
@@ A.3 CO2 Emission Related to Experiments
Experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.057 kgCO2eq/kWh. A cumulative of 23000 hours of computation was performed on hardware of type Tesla V100-SXM2-32GB (TDP of 300W). Total emissions are estimated to be 393.3 kgCO2eq of which 0 percents were directly offset. Estimations were conducted using the MachineLearning Impact calculator333https://mlco2.github.io/impact#compute presented in Lacoste et al., (2019).